OpenStudy (anonymous):

Use the given conditions to write an equation for the line in point-slope form and in slope-intercept form. Passing through (4,7) with x-intercept 6 How do i solve?

OpenStudy (campbell_st):

use the points (4, 7) and (6, 0) to find the gradient m = (7 - 0)/(4 - 6) m = -7/2 use point gradient equation y - 0 = -7/2(x - 6) 2y = 7x + 42

OpenStudy (anonymous):

Thank you